It takes 16.6 min for a 10 ml sample of unknown gas to effuse through a pinhole. A 10 ml sample of helium required 5.0 min. What is the molecular weight of the unknown gas?
1
Expert's answer
2020-11-19T08:51:36-0500

At equal volumes, t2/t1= √M2 / √M1

Let the molecular mass of the unknown gas be x

16.6/5 = √x /4

√x = 16.4 × 4 / 5 = 13.28

x = 13.282

x = 176.36